# Imitating Language via Scalable Inverse Reinforcement Learning

### v1

@@ 1 Introduction
Due to the importance of imitation learning for language modelling, we believe detailed analysis of the underlying imitation problem for sequential decision making is warranted. Despite the widespread perspective of RL for aligning and fine-tuning language models (via RLHF), supervised learning via maximum likelihood estimation for next token prediction remains the dominant component of our imitation learning pipelines due to its simplicity and scalability. However, pure MLE for next token prediction (including “teacher forcing” (Williams and Zipser, 1989)) can create challenges in autoregressive models, many of which being related to classic challenges with behavior cloning (Bain and Sammut, 1995), its equivalent in the context of sequential decision making. Compounding errors can occur due to iterative model application creating data sequences which further shift from the model’s training distribution (Ross et al., 2011; Ji et al., 2023), growing increasingly likely for longer sequences (Dawid and LeCun, 2023). In particular, a model’s own samples can cause such distribution shifts and exposure bias (Ranzato et al., 2016; Bachmann and Nagarajan, 2024; Arora et al., 2022). Taking the RL perspective to imitation aims to mitigate these issues via dynamics-aware optimization, where each action is optimized for the impact on the whole future trajectory. It further enables a shift from passive to active learning, where the system actively generates data. Furthermore, best performance of the fine-tuned model is only one metric of importance. Indeed, continued alignment of language models to human preferences requires sampling for a given prompt, collecting human preferences over these completions, and finally aligning the model via preference fine-tuning. Improving the diversity of sampled completions via temperature sampling Bai et al. (2022), or sampling from a mixture of past models Touvron et al. (2023) has been linked to improvements in downstream performance. Studying inverse RL, and potential divergences, and regularizations provides another angle to increasing diversity (Ghasemipour et al., 2019).
@@
In this paper, we investigate RL-based optimization, in particular the distribution matching perspective to inverse reinforcement learning (IRL), for fine-tuning language models; which can be contrasted with standard MLE as depicted in Figure 1. Our goal is improved understanding of when, and how, IRL can be used as an effective alternative for supervised MLE in the fine-tuning pipeline. The evaluation covers both adversarial and non-adversarial, offline and online methods. We further extend inverse soft Q-learning (Garg et al., 2022a) to explicitly provide a principled connection to classical behavior cloning or MLE. Our experiments range from 250M to 3B parameter models for the encoder-decoder T5 (Raffel et al., 2023) and decoder-only PaLM2 (Anil et al., 2023) models. Throughout evaluation, we investigate both task performance and diversity of model generations illustrating clear benefits of inverse RL over behaviour cloning for imitation learning. A further, principal benefit of RL-centric imitation is the natural connection to later RLHF stages via rewards obtained from demonstration data and we take first steps to analyse the value of IRL-obtained reward functions.
@@
  • We compare MLE and IRL formulations including adversarial and non-adversarial, offline and online methods to improve our understanding of imitation in LLMs. Our main results demonstrate better or on par task performance, with increased diversity of model generations, in particular demonstrating that key improvements can be obtained via (better scalable) offline IRL.
@@ 2 Methods
Language generation can be modeled as sequential decision making problem. On a token level, it is the problem of generating the next token $x_{i}$ given the already generated sequence of tokens $(x_{0},\ldots,x_{i-1})$. We thus seek a distribution $\pi(x_{i}|x_{0},\ldots,x_{i-1})$, which we will also refer to as a policy. For a given policy the likelihood of generating a sequence $\boldsymbol{x}=(x_{0},\ldots,x_{N})$ can be computed autoregressively:
@@
where $D_{f}=E_{\mu_{E}}[f(\frac{\mu_{\pi}(a,s)}{\mu_{E}(a,s)})]$ is an $f$-divergence. Different $f$-divergences have been used in the literature ((Ho and Ermon, 2016; Ghasemipour et al., 2019)).
When taking the example of the reverse KL divergence, with $f(t)=f_{\text{RKL}}(t)=-\log(t)$, we can decompose the objective into a state distribution and an MLE term:
@@
where $\mathrm{KL}$ denotes the KL divergence which corresponds to maximum likelihood (on the discounted state distribution) for the second part after dropping terms independent of $\pi$. In comparison to MLE, IRL algorithms thus also try to match expert actions but additionally attempt to match the state visitations of the expert. The use of different $f$-divergences can influence mode seeking and mode covering properties of the optimal policy (Ghasemipour et al., 2019).
@@
where $r(s,a)$ can be interpreted as an ’implicit’ reward function and we omit its arguments in the following for brevity. Due to the convexity of $f^{*}$ and the concavity of the causal entropy (other terms are linear) this is a saddle point problem (Garg et al., 2022a) and the max-min can be swapped:
@@
Using a telescoping argument (Garg et al., 2022a), we can relate the value of the initial state distribution of a policy to the difference in values on the state-action distribution induced by any arbitrary other policy, i.e. $E_{\rho_{0}}[v^{{r}}]=E_{\mu_{\pi}}[v^{{r}}-\lambda v^{\prime{r}}]$. In this way, we can change the state-action distribution of the second expectation to one induced by an arbitrary policy; in particular we can change it to the one of the expert policy yielding:
@@
With this derivation we show the clear relation between MLE and the inverse RL perspectives: we can see distribution matching (and thus inverse RL) as performing maximum likelihood with a dynamics dependent temporal difference regularization term. In contrast to the adversarial setting, this objective does not require samples from the current policy but uses expert samples only [^2] An online version of IQLearn is derived in Appendix A.1.2.. The regularization term couples the learned policy to a value function and favors policies where the log probability of actions matches the difference in state values. An advantage of the above formulation is that it allows annealing of the regularization term where setting $\lambda=0$ retrieves standard MLE and where by adjusting $\lambda$ the regularization strength can be flexibly increased.
@@ 3.1 Algorithms, baselines, and datasets
In addition to naive maximum likelihood estimation for next token prediction, we evaluate the following IRL methods. Generative adversarial imitation learning (GAIL) (Ho and Ermon, 2016) represents a common adversarial framework training a discriminator to distinguish transitions from agent and dataset and a separate policy. We heuristically adapt the algorithm to mitigate instability for adversarial training (Xing et al., 2021; Wulfmeier et al., 2017) including the start for MLE-trained checkpoints and additional loss terms (including MLE) with further details in Appendix A.1.3. IQLearn (Garg et al., 2022a) departs from adversarial learning and our reformulation from Eq. 14 enables us principled control of the temporal difference regularization component to retain stable training. We will use the reformulated offline variant of the algorithm in all experiments and further add an ablation to its online version in Section 3.3.1. Since we compare all methods with respect to task performance and diversity of generations, we additionally evaluate further entropy regularization terms for the MLE baseline; clearly denoted with ’ent-reg’ in all plots, corresponding to the respective regularization parameter $\lambda$ in GAIL and MLE (see Appendix A.2 for a more detailed description). In line with previous work on inverse RL for language modelling, we apply a short warm-up phase with pure MLE (Cundy and Ermon, 2024; Wu et al., 2021). Rather than separate experiments or heuristic combinations, the explicit MLE term emerging out of our IRL objective in Section 2 enables principled integration of this mechanism.
@@ 3.2.1 T5 models
Figure 4: PaLM2 results for various sampling temperatures with MLE and IQLearn. Left: GSM8k, Mid: TLDR, Right: WMT22, including beam search results. Note, by propagating sequence information during training, IQLearn reduces our inference time dependency on beam search for improving performance.
@@ 3.2.2 PaLM2 models
We also perform experiments with PALM2 models Anil et al. (2023), specifically fine-tuning from a pre-trained PALM2 ’Gecko’ model. We evaluate offline IQLearn on the summarization task TLDR Stiennon et al. (2020), the mathematical reasoning task GSM8k Cobbe et al. (2021), and the large (285M examples) English-to-German translation dataset WMT22 (Kocmi et al., 2022). We limit these experiments to a single choice of the regularization parameter $\lambda$ per task to save computational costs and instead include an analysis of the effect of the sampling temperature parameter during sampling (noting that we similarly observed IQLearn outperforming MLE at varying temperatures for the T5 models). We selected the checkpoints with early stopping for WMT22 as BC was overfitting on the task, unlike TLDR and GSM8K for which we evaluated their latest checkpoints.
@@ 3.3.1 Computational efficiency and accuracy for online & offline inverse RL
One of the key benefits of (offline) MLE compared with (online) IRL-based methods is its lower computational costs. These are principally related to online rollouts - slow, autoregressive sampling from the model. Our re-formulation of IQLearn results in an algorithm that can be applied offline on a fixed dataset, which underlies all IQLearn results presented, mitigating this limitation. In this section, we additionally present update and sampling times across algorithms [^3] Computational resources and time are in practice highly dependent on hardware and implementation but always include crucial, additional sampling costs. and the comparison with the online application of IQLearn (see Appendix A.1.2). Figure 5 demonstrates minimal task performance gains (for the T5 base mode on GSM8k), though considerably improved diversity for model generations. At the same time, Table 1 visualizes the relative cost of sampling in comparison to different algorithm updates, excluding sampling. Note that MLE batch sizes are smaller than IRL ones as we add additional online samples to the batch. Experiment times for online IRL can be reduced, but at the cost of additional hardware, via distributed training with separate sampling infrastructure. The application choice of online or offline IRL finally lies with the practitioner trading of additional computational cost with diversity benefits.
We provide further intuition for the lower differences between offline and online Inverse RL via toy experiments in Appendix A.3.3. At its core, the specific structure of autoregressive language generation, in particular the concatenation underlying single-turn dynamics, prevents exact recovery after mistakes, which could be otherwise learned via IRL in the online setting. Therefore, extensions of the LLM action space to enable recovery can be a fruitful direction for increased benefits from online data with RL and IRL style methods (Cundy and Ermon, 2024).
Table 1: Algorithm profiling with computation times in milliseconds. Sampling refers to sampling a number of sequences equivalent to batch size and often uses equal or more time than updating. These times generally depend on hardware, implementation and code optimization.
@@ 4 Related Work
Understanding language modeling as an imitation problem has been previously explored. Indeed, the link can already be made from MLE, commonly referred to as Behavioral Cloning (BC) Bain and Sammut (1995) from an imitation perspective. Although the link to imitation has been made explicit recently, either theoretically Sun (2024), or in the case of distillation Agarwal et al. (2024); Hubert et al. (2023); Lin et al. (2020); Hormann and Sokolov (2021), some works had already tackled the issues of MLE with imitation-like techniques. For example, adversarial training of text generation an alternative to MLE was first proposed in SeqGAN Yu et al. (2017), and followed-up by a series of work using GANs for text generation Ke et al. (2019). These methods have been shown to work only in the temperature 1 regime Caccia et al. (2018), a possible shortcoming that we address in Section 3. Then, leveraging the literature of imitation learning, GAIL was successfully adapted to language Wu et al. (2021), showing an improvement over MLE. Closer to our contributions, IQLearn was also utilized for language in SequenceMatch Cundy and Ermon (2024). Key differences to our work include the reformulation as temporal difference regularized MLE, comparison with other inverse RL methods and focus on computational costs via the application of offline IQLearn. Indeed, SequenceMatch requires the use of online data, via the introduction of the "backward" token, that allows the model to change a previously chosen token during sampling.
@@ 5 Discussions
The field of imitation learning has led to a gamut of algorithms, many of which are intuitively simple to implement with existing RL or RLHF infrastructure (e.g. (Reddy et al., 2019; Watson et al., 2023; Hejna and Sadigh, 2023)). Ease of adaptation and hyperparameter tuning have principal impact on our practical algorithm choices and the methods and extensions discussed in this work enabled quick first results and iteration. Looking forward, the evaluation and adaptation of further imitation learning methods to LLMs is likely to lead to fruitful results in the coming years. While our analysis focuses on specific algorithms, our key arguments should be seen in the light of the benefits of underlying mechanisms rather than the specific methods.
The sampling-free application of RL mechanism can eventually extend to even larger datasets such as pretraining data, domains with high requirements for computational efficiency. While these datasets are, in a certain sense, less optimal and representative of preferred model behaviour, different aspects of our analysis have the potential to generalize, such as increased efficiency for modelling sequential decision making or increased diversity of model generations. Pretraining and other sub-optimal data could further be used as additional data source for IQLearn and related algorithms without autoregressive sampling.
Finally, RLHF’s key role lies in the alignment of models with respect to user preferences. By integrating SFT data into RL-based optimization, we hope to expand data used for preference description from paired comparisons to individual demonstrations. Integrating generative and discriminative information from different data sources into a unified RL-based framework for unified LLM post-training has further practical potential as previously demonstrated for the generative side (Ouyang et al., 2022). Our reward analysis in Section 3 provides an initial signal about IRL-extracted reward functions including crucial information about LLM task performance.

### v2

@@ 1 Introduction
Due to the importance of imitation for language modelling, we believe detailed analysis of the underlying problem for sequential decision making is warranted. Despite the widespread perspective of RL for aligning and fine-tuning language models (via RLHF), supervised learning via maximum likelihood estimation for next token prediction remains the dominant component of our imitation learning pipelines due to its simplicity and scalability. However, pure MLE for next token prediction (including “teacher forcing” (Williams and Zipser, 1989)) can create challenges in autoregressive models, many of which being related to classic challenges with behavior cloning (Bain and Sammut, 1995), its equivalent in the context of sequential decision making. Compounding errors can occur due to iterative model application creating data sequences which further shift from the training distribution (Ross et al., 2011; Ji et al., 2023), growing increasingly likely for longer sequences (Dawid and LeCun, 2023). In particular, a model’s own samples can cause such distribution shifts and exposure bias (Ranzato et al., 2016; Bachmann and Nagarajan, 2024; Arora et al., 2022). Taking the RL perspective to imitation aims to mitigate these issues via dynamics-aware optimization, where each action is optimized for the impact on the whole future trajectory. It further enables a shift from passive to active learning, where the system actively generates data. Furthermore, best performance of the fine-tuned model is only one metric of importance. Indeed, continued alignment of language models to human preferences requires sampling for a given prompt, collecting human preferences over these completions, and finally aligning the model via preference fine-tuning. Improving the diversity of sampled completions via temperature sampling Bai et al. (2022), or using mixtures of past models Touvron et al. (2023) has been linked to performance improvements. Studying inverse RL, potential divergences, and regularizations provides another angle to increasing diversity (Ghasemipour et al., 2019).
@@
In this paper, we investigate RL-based optimization, in particular the distribution matching perspective to inverse reinforcement learning (IRL), for fine-tuning language models; which can be contrasted with standard MLE as depicted in Figure 1. Our goal is improved understanding of when, and how, IRL can be used as an effective alternative for supervised MLE in the fine-tuning pipeline. The evaluation covers both adversarial and non-adversarial, offline and online methods. We further extend inverse soft Q-learning (Garg et al., 2022a) to explicitly provide a principled connection to classical behavior cloning or MLE. Our experiments range from 250M to 3B parameter models for the encoder-decoder T5 (Raffel et al., 2023) and decoder-only PaLM2 (Anil et al., 2023) models. Throughout evaluation, we investigate task performance and diversity of model generations illustrating clear benefits of inverse RL over behavior cloning for imitation. A further, principal benefit of RL-centric imitation – close to RLAIF, RL from Artificial Intelligence Feedback – is the natural connection to later RLHF stages via rewards obtained from demonstration data and we take first steps to analyse the value of IRL-obtained reward functions.
@@
  • We compare MLE and IRL formulations including adversarial and non-adversarial, offline and online methods to improve our understanding of imitation in LLMs. Our main results show better or on par task performance, with increased diversity of model generations, in particular demonstrating that key improvements can be obtained via (better scalable) offline IRL.
@@ 2 Methods
Language generation can be modeled as a sequential decision making problem. On a token level, it is the problem of generating the next token $x_{i}$ given the already generated sequence of tokens $(x_{0},\ldots,x_{i-1})$. We thus seek a distribution $\pi(x_{i}|x_{0},\ldots,x_{i-1})$, which we will also refer to as a policy. For a given policy the sequence likelihood $\boldsymbol{x}=(x_{0},\ldots,x_{N})$ can be computed autoregressively:
@@
where $D_{f}=E_{\mu_{E}}[f(\frac{\mu_{\pi}(a,s)}{\mu_{E}(a,s)})]$ is an $f$-divergence. Different $f$-divergences have been used in the literature (Ho and Ermon, 2016; Ghasemipour et al., 2019). When taking the example of the reverse KL divergence, with $f(t)=f_{\text{RKL}}(t)=-\log(t)$, we can decompose the objective into a state distribution and an MLE term:
@@
where $\mathrm{KL}$ denotes the KL-divergence which corresponds to maximum likelihood (on the discounted state distribution) for the second part after dropping terms independent of $\pi$. In comparison to MLE, IRL algorithms thus also try to match expert actions but additionally attempt to match the state visitations of the expert. The use of different $f$-divergences can influence mode seeking and mode covering properties of the optimal policy (Ghasemipour et al., 2019).
@@
where $r(s,a)$ can be interpreted as an ‘implicit’ reward function and we omit its arguments in the following for brevity. Due to the convexity of $f^{*}$ and the concavity of the causal entropy (other terms are linear) this is a saddle point problem (Garg et al., 2022a) and the max-min can be swapped:
@@
Using a telescoping argument (Garg et al., 2022a; Sikchi et al., 2023), we can relate the value of the initial state distribution of a policy to the difference in values on the state-action distribution induced by arbitrary other policies, i.e. $E_{\rho_{0}}[v^{{r}}]=E_{\mu_{\pi}}[v^{{r}}-\lambda v^{\prime{r}}]$. In particular, here we can change the second expectation’s state-action distribution to one induced by the expert policy yielding:
@@
With this derivation we show the clear relation between MLE and the inverse RL perspectives (expanding on (Piot et al., 2014)): we can see distribution matching (and thus inverse RL) as performing maximum likelihood with a dynamics dependent temporal difference regularization term. In contrast to the adversarial setting, this objective does not require samples from the current policy but uses expert samples only [^2] An online version of IQLearn is derived in Appendix A.1.2.. The regularization term couples the learned policy to a value function and favors policies where the log probability of actions matches the difference in state values. An advantage of the above formulation is that it allows annealing of the regularization term where setting $\lambda=0$ retrieves standard MLE and where by adjusting $\lambda$ the regularization strength can be flexibly increased.
@@ 3.1 Algorithms, baselines, and datasets
In addition to naive maximum likelihood estimation for next token prediction, we evaluate the following IRL methods. Generative adversarial imitation learning (GAIL) (Ho and Ermon, 2016) represents a common adversarial framework training a discriminator to distinguish transitions from agent and dataset and a separate policy. We heuristically adapt the algorithm to mitigate instability for adversarial training (Xing et al., 2021; Wulfmeier et al., 2017) including the start for MLE-trained checkpoints and additional loss terms (including MLE) with further details in Appendix A.1.3. IQLearn (Garg et al., 2022a) departs from adversarial learning and our reformulation from Eq. 14 enables us principled control of the temporal difference regularization component to retain stable training. We will use the reformulated offline variant of the algorithm in all experiments and further add an ablation to its online version in Section 3.3.1. Since we compare all methods with respect to task performance and diversity of generations, we additionally evaluate further entropy regularization terms for the MLE baseline; clearly denoted with ‘ent-reg’ in all plots, corresponding to the respective regularization parameter $\lambda$ in GAIL and MLE (see Appendix A.2 for a more detailed description). In line with previous work on inverse RL for language modelling, we apply a short warm-up phase with pure MLE (Cundy and Ermon, 2024; Wu et al., 2021). Rather than separate experiments or heuristic combinations, the explicit MLE term emerging out of our IRL objective in Section 2 enables principled integration of this mechanism.
@@ 3.2.1 T5 models
Figure 4: PaLM2 results for various sampling temperatures with MLE and IQLearn. Left: GSM8k, Mid: TLDR, Right: WMT22, including beam search. By propagating sequence information during training, IQLearn reduces inference time dependency on beam search for improving performance.
@@ 3.2.2 PaLM2 models
We also perform experiments with PALM2 models Anil et al. (2023), specifically fine-tuning from a pre-trained PALM2 ‘Gecko’ model. We evaluate offline IQLearn on the summarization task TLDR Stiennon et al. (2020), the mathematical reasoning task GSM8k Cobbe et al. (2021), and the large (285M examples) English-to-German translation dataset WMT22 (Kocmi et al., 2022). We limit these experiments to a single choice of the regularization parameter $\lambda$ per task to save computational costs and instead include an analysis of the effect of the sampling temperature parameter during sampling (noting that we similarly observed IQLearn outperforming MLE at varying temperatures for the T5 models). We selected the checkpoints with early stopping for WMT22 as BC was overfitting on the task, unlike TLDR and GSM8K for which we evaluated their latest checkpoints.
@@ 3.3.1 Computational efficiency and accuracy for online & offline inverse RL
One of the key benefits of (offline) MLE compared with (online) IRL-based methods is its lower computational costs. These are principally related to online rollouts – slow, autoregressive sampling from the model. Our re-formulation of IQLearn results in an algorithm that can be applied offline on a fixed dataset, which underlies all IQLearn results presented, mitigating this limitation. In this section, we additionally present update and sampling times across algorithms [^3] Computational resources and time are in practice highly dependent on hardware and implementation but always include crucial, additional sampling costs. and the comparison with the online application of IQLearn (see Appendix A.1.2). Figure 5 demonstrates minimal task performance gains (for the T5 base mode on GSM8k), though considerably improved diversity for model generations. At the same time, Table 1 visualizes the relative cost of sampling in comparison to different algorithm updates, excluding sampling. Note that MLE batch sizes are smaller than IRL ones as we add additional online samples to the batch. Experiment times for online IRL can be reduced, but at the cost of additional hardware, via distributed training with separate sampling infrastructure. The application choice of online or offline IRL finally lies with the practitioner trading of additional computational cost with diversity benefits.
We provide intuition for low differences between offline and online Inverse RL via toy experiments in Appendix A.3.3. At its core, the dynamics of autoregressive language generation, in particular the concatenation underlying single-turn settings, prevents exact recovery after mistakes, which could be otherwise learned via online IRL. Therefore, extensions of the LLM action space to enable recovery is a fruitful direction for increased benefits from online data with RL or IRL methods (Cundy and Ermon, 2024).
Table 1: Algorithm profiling with computation in milliseconds. ‘Sampling’ refers to generating a number of sequences equivalent to batch size and often uses equal or more time than updates. These times depend on hardware, implementation and code optimization.
@@ 4 Related Work
Understanding language modeling as an imitation problem has been previously explored. Indeed, the link can already be made from MLE, commonly referred to as Behavioral Cloning (BC) Bain and Sammut (1995) from an imitation perspective. Although the link to imitation has been made explicit recently, either theoretically Sun (2024), or in the case of distillation and divergence minimisation Agarwal et al. (2024); Hubert et al. (2023); Lin et al. (2020); Hormann and Sokolov (2021); Li et al. (2024b), some works had already tackled the issues of MLE with imitation-like techniques. For example, adversarial training of text generation an alternative to MLE was first proposed in SeqGAN Yu et al. (2017), and followed-up by a series of work using GANs for text generation Ke et al. (2019). These methods have been shown to work only in the temperature 1 regime Caccia et al. (2018), a possible shortcoming that we address in Section 3. Then, leveraging the literature of imitation learning, GAIL was successfully adapted to language Wu et al. (2021), showing an improvement over MLE. Closer to our contributions, IQLearn is utilized for language in SequenceMatch Cundy and Ermon (2024). Key differences to our work include the reformulation as temporal difference regularized MLE, large-scale analysis including other inverse RL methods and focus on computational costs via the application of offline IQLearn. Indeed, SequenceMatch requires the use of online data, via the introduction of the "backward" token, that allows the model to change a previously chosen token during sampling.
@@ 5 Discussions
The field of imitation learning has led to a gamut of algorithms, many of which are intuitively simple to implement with existing RL or RLHF infrastructure (e.g. (Reddy et al., 2019; Watson et al., 2023; Hejna and Sadigh, 2023; Rafailov et al., 2024)). Ease of adaptation and hyperparameter tuning have principal impact on our practical algorithm choices and the methods and extensions discussed in this work enabled quick first results and iteration. Looking forward, the evaluation and adaptation of further imitation learning methods to LLMs is likely to lead to fruitful results in the coming years. While our analysis focuses on specific algorithms, our key arguments should be seen in the light of the benefits of underlying mechanisms rather than the specific methods.
The sampling-free application of RL mechanism can eventually extend to even larger datasets such as pretraining data, domains with high requirements for computational efficiency. While these datasets are, in a certain sense, less optimal and representative of preferred model behavior, different aspects of our analysis have the potential to generalize, such as increased efficiency for modelling sequential decision making or increased diversity of model generations. Pretraining and other sub-optimal data could further be used as additional data source for IQLearn and related algorithms (Ma et al., 2022; Kim et al., 2022; Sikchi et al., 2023) instead of autoregressive sampling data.
Finally, RLHF’s key role lies in the alignment of models with respect to user preferences. By integrating SFT data into RL-based optimization, we hope to expand data used for preference description from paired comparisons to individual demonstrations. Integrating generative and discriminative information from different data sources into a unified RL-based framework for unified LLM post-training has further practical potential as previously demonstrated for the generative side (Ouyang et al., 2022) with first steps towards reward transfer (Li et al., 2024a). Our reward analysis in Section 3 provides an initial signal about IRL-extracted reward functions including crucial information about LLM task performance.